# Hertzsprung gap stars in nearby galaxies and the Quest for Luminous Red Novae Progenitorsthanks: Tables 3 and 4 are only available in electronic form at the CDS via anonymous ftp to cdsarc.cds.unistra.fr (130.79.128.5) or via https://cdsarc.cds.unistra.fr/cgi-bin/qcat?J/A+A/

### v1

@@ Abstract
Results. Our sample includes 246,573 YSG candidates with masses between $3M_{\odot}$ and $20M_{\odot}$ and is affected by various contaminants, notably foreground stars and extinguished main-sequence stars. After excluding foreground stars using Gaia proper motions, contamination is estimated at 1.7% from foreground stars (based on TRILEGAL simulations) and $\sim$20from extinction affecting main-sequence stars. Combining our YSG candidates with time-domain catalogues yielded several interesting candidates. In particular, we identified twelve LRN precursor candidates for which followup is encouraged.
@@ 2.1 Selection criteria
To get a complete sample of nearby galaxies, we use the HECATE catalogue of galaxies (Heraklion Extragalactic CATaloguE, Kovlakas et al. 2021), based on the HyperLEDA database (Paturel et al., 2003) but supplemented with robust distance estimates, galaxy mass, metallicity and star formation rate. The base sample, containing 1883 galaxies closer than 20 Mpc, is crossmatched with HST observations available on the Mikulski Archive for Space Telescopes (MAST), to find galaxies with overlapping deep ($t_{\mathrm{exp}}>300$ s) exposures in two optical filters [^2] This crossmatch was performed using the API of MAST, available as part of the astroquery v0.4.6 package (Ginsburg et al., 2019). The list of selected filters is detailed in Table 1, together with the number of galaxies selected for each filter. We select the $F475W$, $F555W$, $F606W$ and $F814W$ filters because they have responses close to the Johnson-Cousins $V$ ($F475W$, $F555W$, $F606W$) and $I$ ($F814W$) filters, allowing us to study $V$ against $V-I$ colour-magnitude diagrams (CMD). The resulting HST-observed sample contains 575 galaxies. The distributions of distance and angular sizes are shown in the top panel of Figure 1.
@@ 2.2 HST data retrieval
The MAST third party catalogues were produced at the time of the data reduction of the corresponding observations, using either DAOPHOT (Stetson, 1987), Sextractor (Bertin & Arnouts, 1996), or the Hubble Advanced Products pipeline (Tran et al., 2020). For both types of catalogues, magnitudes are retrieved in the large aperture photometry (MagAper2). We made sure that MAST catalogues are reliable and have well calibrated photometry, by comparing their sky coverage and their CMD with HSCv3 data for two test galaxies (NGC 45 and ESO 209-9). In all tested filters, the bias and standard deviation between HSCv3 and MAST photometry is $<0.05$ mag and $<0.2$ mag, respectively. Once downloaded, all catalogues of a given galaxy are merged and crossmatched to obtain a single “master” catalogue, using astropy v5.3.3 (Astropy Collaboration et al., 2013) [^4] A data-dependent matching radius $r_{1}$ ($20.3<\frac{r_{1}}{\mathrm{arcsec}}<$ 2) is used to associate sources from different filters, and another one $r_{2}$ ($0.30.1<\frac{r_{2}}{\mathrm{arcsec}}<$ 0.3) is used to eliminate duplicate sources.. To discard compact star clusters and cosmic rays, we discarded sources with concentration indexes (CI) $<0.85$ or magnitude errors $>0.1$.
@@
Table 1: HST filters used in this study. The last column gives the number of galaxies selected in the corresponding filters.
@@ 2.3.1 Stellar evolution tracks
To analyse the CMD of galaxies, a physical reference is needed, such as isochrones or stellar evolution tracks in the observer’s frame. However, the shape of an observed CMD heavily depends on intrinsic and extrinsic parameters. Intrinsic parameters include the galaxy star formation history and its metallicity, while extrinsic parameters correspond to measurement uncertainties and the extinction along the line of sight, commonly modelled with the extinction parameter $A_{\lambda}$. Using the synthetic photometry of stellar evolution tracks provided by MIST (MESA Isochrones & Stellar Tracks [^5] https://waps.cfa.harvard.edu/MIST, Dotter 2016; Choi et al. 2016), we generated tracks for a grid of zero-age main sequence (ZAMS) stellar masses, extinctions, and metallicities, summarised in Table 2. The ranges of extinctions and metallicities correspond to the ranges of estimated Galactic extinctions and metallicities of galaxies in our sample, and the mass range encompasses most of known LRNe progenitors (e.g. MacLeod et al. 2022). The upper limit of the mass range does not restrict the YSG selection; instead, it marks the point beyond which stellar masses become poorly estimated, as the Hertzsprung gap can no longer be consistently defined as a continuous post-MS phase.
@@ 2.4.1 Removing known contaminants
As an initial cleaning process, we crossmatched between our sample and the Simbad [^8] https://simbad.cds.unistra.fr/ and Gaia DR3 (Gaia Collaboration et al., 2023b) catalogues using a matching radius of 0.5”. Simbad quasars, globular clusters, and classical cepheids were discarded. Additional quasars were found and excluded using the Milliquas (Flesch, 2021) and Gaia Extragalactic (Gaia Collaboration et al., 2023a) catalogues. However, we flagged known LBVs and yellow/red supergiants for further consideration within the sample. Sources exhibiting significant Gaia proper motion ($\mathrm{PM}/\mathrm{PM_{err}}>4$) are identified as Galactic and discarded. Although other types of contaminants cannot be directly excluded, their presence can be quantified statistically.
@@ 2.4.3 Extinguished main-sequence stars
Although we account for the Galactic extinction along the line of sight for each galaxy using the appropriate MIST model, intrinsic extinction inside each galaxy is not modelled at this stage. Consequently, a significant fraction of Hertzsprung gap sources may actually be extinguished MS stars. To quantify this, we assigned to each candidate a probability to be an extinguished main-sequence star. Assuming a normal distribution for extinction, we measured extinction scatter by evaluating the standard deviation of $A_{V}$ in a thin region of the CMD, chosen to be the red giant branch, as illustrated in Figure 5. This standard deviation is used to model the additional extinction in the selection region, resulting in a probability map for each CMD (Figure 5, bottom panel). This probability is naturally highest at the blue edge of the selection region, which separates the main-sequence region from the Hertzsprung gap. By averaging these probabilities on a per-galaxy basis, we can determine the contamination rate due to extinction. The distribution of $A_{V}$ scatters estimated for all selected galaxies is shown in Figure 6. Besides providing contamination probabilities, it also suggests that the 0.4 mag extinction applied to the red edge of the selection region is sufficient to recover most of the extinguished YSG (Section 2.3.2).
Figure 5: (Top) Density plot of the RGB of IC 1613 in the (Magnitude - $A_{V}$) plane. The top curve shows the KDE of $A_{V}$ values, with a standard deviation of $\pm$0.21 (dotted lines). (Bottom) Resulting probability map to be an extinguished MS star in the selection region of YSG sources in the IC 1613 galaxy.
@@ 3.1 Selected candidates
Our analysis identified a total of 246,573 YSG candidates, a portion of which being listed in Table 3. The cumulative distribution of their distances is depicted in Figure 7, indicating that about 25% of these candidates are located in galaxies at approximately 0.8 Mpc. These candidates are actually located in M 31 (53,323 candidates) and M 33 (15,706), the two galaxies with most candidates. The three next galaxies with the most YSG candidates are M101 (15,156), NGC 300 (9,706), and NGC 253 (8,004).
As a result of the CMD analysis, we identified YSG candidates in 355 galaxies. The HST data provenance, available blue/green filter, coverage, and the parameters used for fitting the CMD as described in Section 2.3 are provided for each galaxy in Table 4. This table also details the number of candidates for each galaxy and the estimated contribution from different contaminants.
@@
1 Dra dSph 260.47560 57.95323 16.8 3.8 3.75 0.07
2 Dra dSph 260.10617 57.97408 15.6 6.0 3.73 0.00
3 Dra dSph 260.10771 57.97393 16.3 4.3 3.79 0.17
4 Dra dSph 260.07129 57.97381 15.9 5.4 3.75 0.01
5 Dra dSph 260.09113 57.97427 16.9 3.8 3.84 0.42
6 Dra dSph 260.20050 57.91569 17.6 3.2 3.74 0.02
7 Dra dSph 260.10907 57.97411 15.5 5.9 3.73 0.01
8 Dra dSph 260.07055 57.97380 15.7 5.7 3.74 0.00
9 Dra dSph 260.06858 57.97374 15.8 5.6 3.72 0.00
10 Dra dSph 260.09112 57.97362 17.0 3.6 3.78 0.13
11 Dra dSph 260.10544 57.97412 15.7 5.6 3.74 0.01
12 ESO 104-22 283.90293 -64.80823 22.2 15.7 3.63 0.00
13 ESO 104-22 283.92116 -64.80982 23.2 11.2 3.85 0.75
14 ESO 104-22 283.91606 -64.80495 21.9 15.7 3.67 0.02
15 ESO 104-22 283.91053 -64.80493 24.0 8.8 3.69 0.04
16 ESO 104-22 283.93422 -64.81086 22.6 13.9 3.64 0.00
17 ESO 104-22 283.93325 -64.80926 22.7 12.2 3.73 0.12
18 ESO 104-22 283.90344 -64.80865 22.4 14.0 3.79 0.37
19 ESO 104-22 283.93478 -64.81320 24.0 8.7 3.69 0.04
20 ESO 104-22 283.92837 -64.81673 24.0 8.8 3.69 0.04
21 ESO 104-22 283.90688 -64.81096 22.4 13.5 3.77 0.27
22 ESO 104-22 283.92554 -64.81018 24.1 9.0 3.82 0.51
23 ESO 104-44 287.85375 -64.22362 24.3 8.9 3.67 0.01
24 ESO 104-44 287.85099 -64.22005 23.7 10.8 3.65 0.00
25 ESO 104-44 287.85434 -64.22405 24.2 8.8 3.82 0.44
26 ESO 104-44 287.85947 -64.22657 24.4 8.8 3.90 0.97
27 ESO 104-44 287.85444 -64.21808 21.9 15.4 3.75 0.12
28 ESO 104-44 287.85261 -64.22444 23.8 10.2 3.65 0.00
29 ESO 104-44 287.85741 -64.21309 22.2 14.2 3.63 0.00
Table 3: List of YSG candidates. Column 2: host galaxy. Columns 3–5: HST coordinates (deg) and magnitude (mag). Columns 6–7: MIST inferred stellar parameters. Column 8: Probability to be an extinguished MS star. The complete list will be made available in electronic form.
Galaxy Distance HSTcov Source Camera / Filter Nsrc MIST M∗,ref $A_{V}$ [Fe/H] $n_{\mathrm{GM}}$ Ncand $f_{\mathrm{extin}}$ $f_{\mathrm{foreground}}$
@@
M 31 0.8 34.3 HSCv3+MAST ACS $F475W$ 97929 4 0.49 -0.88 0 53323 0.251 0.136
M 33 0.8 10.0 HSCv3 ACS $F475W$ 2834135 4 0.51 -1.13 10 15706 0.183 0.137
Sgr dIG 1.1 100.0 HSCv3 ACS $F606W$ 13513 3 0.20 -1.47 50 357 0.306 0.446
Sex A 1.4 72.9 HSCv3 WFPC2 $F555W$ 5710 4 0.16 -1.19 20 499 0.185 0.028
NGC 55 2.2 10.7 HSCv3 ACS $F606W$ 146004 6 0.13 -1.45 10 2630 0.352 0.003
NGC 7793 3.4 57.0 HSCv3 ACS $F555W$ 90987 8 0.06 -0.57 50 946 0.226 0.029
NGC 5253 3.5 100.0 HSCv3 ACS $F555W$ 62144 8 0.11 -1.02 50 801 0.239 0.046
NGC 247 3.8 35.5 HSCv3 ACS $F606W$ 277045 8 0.03 -0.55 20 512 0.294 0.045
NGC 253 3.9 41.5 HSCv3 ACS $F606W$ 634859 8 0.37 -0.93 10 8004 0.353 0.005
ESO 223-9 4.4 100.0 HSCv3 ACS $F606W$ 15109 8 0.60 -1.05 90 486 0.267 0.516
ESO 137-18 4.7 96.7 HSCv3 ACS $F606W$ 19885 8 0.54 -0.83 30 578 0.193 0.437
M 83 4.8 37.7 HSCv3 ACS $F555W$ 201354 8 0.16 -0.83 10 7158 0.287 0.031
NGC 45 5.8 82.3 HSCv3 ACS $F555W$ 13503 8 0.19 -0.61 50 470 0.190 0.035
IC 5052 5.9 75.8 HSCv3 ACS $F606W$ 39959 8 0.17 -0.98 50 929 0.332 0.042
IC 1959 6.4 100.0 HSCv3 ACS $F606W$ 12641 8 0.09 -0.86 50 302 0.333 0.007
NGC 1566 6.5 47.3 HSCv3 ACS $F555W$ 21651 8 -0.01 -0.63 20 2762 0.221 0.012
Fourcade-Figueroa 6.7 48.1 HSCv3 ACS $F606W$ 38012 8 0.13 -1.26 0 593 0.288 0.136
NGC 3621 7.2 45.1 HSCv3 ACS $F555W$ 67001 8 0.17 -1.01 0 6950 0.288 0.012
M 74 7.3 53.1 HSCv3 ACS $F555W$ 59479 8 0.38 -0.86 20 6548 0.229 0.007
IC 4710 7.4 99.9 HSCv3 ACS $F606W$ 26877 8 0.18 -0.73 20 920 0.280 0.126
UGC 3755 7.4 100.0 HSCv3 ACS $F606W$ 4554 8 0.35 -1.03 50 301 0.236 0.021
NGC 2188 8.0 93.4 HSCv3 ACS $F606W$ 16356 8 0.20 -0.63 50 1821 0.271 0.010
NGC 1433 8.3 58.9 HSCv3 WFC3 $F555W$ 2308 10 0.14 -0.47 50 329 0.201 0.046
IC 4870 8.5 100.0 HSCv3 ACS $F606W$ 4163 8 0.38 -0.97 80 532 0.186 0.015
NGC 7090 8.6 60.4 HSCv3 ACS $F606W$ 14292 10 0.51 -1.12 50 1091 0.375 0.014
NGC 24 9.1 93.1 HSCv3 ACS $F606W$ 18596 8 0.18 -1.26 0 881 0.366 0.009
NGC 4517 9.7 20.2 HSCv3 ACS $F606W$ 3632 8 0.07 -0.50 70 314 0.155 0.023
ESO 373-8 9.8 71.9 HSCv3 ACS $F606W$ 5719 10 0.51 -0.97 0 377 0.302 0.102
M 95 10.0 41.0 HSCv3 WFC3 $F555W$ 3101 10 0.12 -0.98 0 405 0.211 0.015
IC 5201 10.3 72.8 HSCv3 ACS $F606W$ 22810 10 0.14 -0.48 50 729 0.231 0.045
M 66 10.4 36.7 HSCv3 WFC3 $F555W$ 29236 12 0.15 -0.56 0 2423 0.137 0.004
NGC 4592 11.2 98.5 HSCv3 WFC3 $F555W$ 9412 10 -0.01 -0.88 10 828 0.208 0.007
NGC 5398 11.4 99.8 HSCv3 ACS $F606W$ 14604 8 0.33 -1.11 0 1342 0.391 0.028
NGC 3810 16.2 94.4 HSCv3 ACS $F555W$ 7386 16 0.40 -0.93 50 394 0.182 0.005
@@
Table 4: List of galaxies having more than 100 YSG candidates, sorted by distance. Column 2: distance to the galaxy. Columns 3–6: information on their HST dataset. Columns 7–10: parameters used to fit their CMD. Column 11: number of resulting candidates. Columns 12–13: contamination estimates. The complete list of galaxies having YSG candidate will be made available in electronic form.

### v2

@@ Abstract
Results. Our sample includes 154,494 YSG candidates with masses between $3M_{\odot}$ and $20M_{\odot}$ and is affected by various contaminants, notably foreground stars and extinguished main-sequence stars. After excluding foreground stars using Gaia proper motions, contamination is estimated at 1% from foreground stars (based on TRILEGAL simulations) and $\sim$20from extinction affecting main-sequence stars. Combining our YSG candidates with time-domain catalogues yielded several interesting candidates. In particular, we identified twelve LRN precursor candidates for which followup is encouraged.
@@ 2.1 Selection criteria
To get a complete sample of nearby galaxies, we use the HECATE catalogue of galaxies (Heraklion Extragalactic CATaloguE, Kovlakas et al. 2021), based on the HyperLEDA database (Paturel et al., 2003) but supplemented with robust distance estimates, galaxy mass, metallicity and star formation rate. The base sample, containing 1883 galaxies closer than 20 Mpc, is crossmatched with HST observations available on the Mikulski Archive for Space Telescopes (MAST), to find galaxies with overlapping deep ($t_{\mathrm{exp}}>300$ s) exposures in two optical filters [^2] This crossmatch was performed using the API of MAST, available as part of the astroquery v0.4.6 package (Ginsburg et al., 2019).. Only galaxies having at least 2% of their area covered by HST are considered, discarding de facto the Magellanic Clouds [^3] Recent studies targeting the Magellanic Clouds using ground-based surveys and Swift-UVOT provide YSG samples of good completeness (Neugent et al., 2010, 2012; Humphreys et al., 2023; O’Grady et al., 2024).. The list of selected filters is detailed in Table 1, together with the number of galaxies selected for each filter. We select the $F475W$, $F555W$, $F606W$ and $F814W$ filters because they have responses close to the Johnson-Cousins $V$ ($F475W$, $F555W$, $F606W$) and $I$ ($F814W$) filters, allowing us to study $V$ against $V-I$ colour-magnitude diagrams (CMD). The resulting HST-observed sample contains 575 galaxies. The distributions of distance and angular sizes are shown in the top panel of Figure 1.
@@ 2.2 HST data retrieval
The MAST third party catalogues were produced at the time of the data reduction of the corresponding observations, using either DAOPHOT (Stetson, 1987), Sextractor (Bertin & Arnouts, 1996), or the Hubble Advanced Products pipeline (Tran et al., 2020). For both types of catalogues, magnitudes are retrieved in the large aperture photometry (MagAper2). We made sure that MAST catalogues are reliable and have well calibrated photometry, by comparing their sky coverage and their CMD with HSCv3 data for two test galaxies (NGC 45 and ESO 209-9). In all tested filters, the bias and standard deviation between HSCv3 and MAST photometry is $<0.05$ mag and $<0.2$ mag, respectively. Once downloaded, all catalogues of a given galaxy are merged and crossmatched to obtain a single “master” catalogue, using astropy v5.3.3 (Astropy Collaboration et al., 2013) [^5] A data-dependent matching radius $r_{1}$ ($20.3<\frac{r_{1}}{\mathrm{arcsec}}<$ 2) is used to associate sources from different filters, and another one $r_{2}$ ($0.30.1<\frac{r_{2}}{\mathrm{arcsec}}<$ 0.3) is used to eliminate duplicate sources. $r_{2}$ is the local minimum of the distribution of separations in the interval [0.1,0.3] arcsec. This choice allows us to eliminate duplicate sources while keeping close neighbors, given the different levels of astrometric accuracies when using different source-extracting algorithms (sextractor, daophot).. To discard compact star clusters, cosmic rays and sources affected by confusion, we discarded sources with concentration indexes (CI) $<0.85$ or $>1.5$ or magnitude errors $>0.1$.
@@
 [^6] HST filters used in this study. The last column gives the number of galaxies selected in the corresponding filters.
@@ 2.3.1 Stellar evolution tracks
To analyse the CMD of galaxies, a physical reference is needed, such as isochrones or stellar evolution tracks in the observer’s frame. However, the shape of an observed CMD heavily depends on intrinsic and extrinsic parameters. Intrinsic parameters include the galaxy star formation history and its metallicity, while extrinsic parameters correspond to measurement uncertainties and the extinction along the line of sight, commonly modelled with the extinction parameter $A_{\lambda}$. Using the synthetic photometry of stellar evolution tracks provided by MIST (MESA Isochrones & Stellar Tracks [^7] https://waps.cfa.harvard.edu/MIST, Dotter 2016; Choi et al. 2016), we generated tracks for a grid of zero-age main sequence (ZAMS) stellar masses, extinctions, and metallicities, summarised in Table 2. The ranges of extinctions and metallicities correspond to the ranges of estimated Galactic extinctions and metallicities of galaxies in our sample, and the mass range encompasses most of known LRNe progenitors (e.g. MacLeod et al. 2022). The upper limit of the mass range does not restrict the YSG selection; instead, it marks the point beyond which stellar masses become poorly estimated, as the Hertzsprung gap can no longer be consistently defined as a continuous post-MS phase. It is important to note that our selection will contain YSG regardless of their stellar evolution phase, either post-MS or post-RSG. While some YSG have been identified as post-RSG in recent studies (e.g. Humphreys et al. 2023), we cannot distinguish between them in our catalog. However, post-RSG YSG typically show only a fast, low-amplitude variability, therefore we expect a stable behaviour in future LSST observations.
Table 2: MIST parameter grid.
@@ 2.4.1 Removing known contaminants
As an initial cleaning process, we crossmatched between our sample and the Simbad [^10] https://simbad.cds.unistra.fr/ and Gaia DR3 (Gaia Collaboration et al., 2023b) catalogues using a matching radius of 0.5”. Simbad quasars, globular clusters, and classical cepheids were discarded. Additional quasars were found and excluded using the Milliquas (Flesch, 2021) and Gaia Extragalactic (Gaia Collaboration et al., 2023a) catalogues. However, we flagged known LBVs, YSGs and red supergiants (RSG) for further consideration within the sample. Sources exhibiting significant Gaia proper motion ($\mathrm{PM}/\mathrm{PM_{err}}>4$) are identified as Galactic and discarded. Although other types of contaminants cannot be directly excluded, their presence can be quantified statistically.
@@ 2.4.3 Extinguished main-sequence stars
Although we account for the Galactic extinction along the line of sight for each galaxy using the appropriate MIST model, intrinsic extinction inside each galaxy is not modelled at this stage. Consequently, a significant fraction of Hertzsprung gap sources may actually be extinguished MS stars. To quantify this, we assigned to each candidate a probability to be an extinguished main-sequence star. Assuming a normal distribution for extinction, we measured extinction scatter by evaluating the standard deviation of $A_{V}$ in a thin region of the CMD, chosen to be the red giant branch, as illustrated in Figure 5. This standard deviation is used to model the additional extinction in the selection region, resulting in a probability map for each CMD (Figure 5, bottom panel). This probability is naturally highest at the blue edge of the selection region, which separates the main-sequence region from the Hertzsprung gap. By averaging these probabilities on a per-galaxy basis, we can determine the contamination rate due to extinction. The distribution of $A_{V}$ scatters estimated for all selected galaxies is shown in Figure 6. Besides providing contamination probabilities, it also suggests that the 0.4 mag extinction applied to the red edge of the selection region is sufficient to recover most of the extinguished YSG (Section 2.3.2). As a result, some RSGs are included in the sample, as discussed in Section 5.1 for the case of M31. However, obscured YSGs and RSGs have also been identified as progenitors of certain infrared transients thought to share similarities with LRNe (Jencson et al., 2019).
Figure 5: (Top) Density plot of the RGB of IC 1613 in the (Magnitude - relative $A_{V}$) plane. The x-axis zeropoint corresponds to the mean of the distribution. The top curve shows the KDE of $A_{V}$ values, with a standard deviation of $\pm$0.21 (dotted lines). (Bottom) Resulting probability map to be an extinguished MS star in the selection region of YSG sources in the IC 1613 galaxy.
@@ 3.1 Selected candidates
Our analysis identified a total of 154,494 YSG candidates before cleaning, a portion of which being listed in Table 3. The cumulative distribution of their distances is depicted in Figure 7, indicating that about 25% of these candidates are located in galaxies at approximately 0.8 Mpc. These candidates are actually located in M31 (40,566 candidates) and M33 (11,603), the two galaxies with most candidates. The three next galaxies with the most YSG candidates are M101 (8,885), NGC 300 (7,525), and NGC 253 (4,000).
As a result of the CMD analysis, we identified YSG candidates in 353 galaxies. The HST data provenance, available blue/green filter, coverage, and the parameters used for fitting the CMD as described in Section 2.3 are provided for each galaxy in Table 4. This table also details the number of candidates for each galaxy and the estimated contribution from different contaminants.
Table 3: List of YSG candidates.
@@
1 WLM Galaxy 0.46607 -15.50511 20.7 6.4 3.70 0.000
2 WLM Galaxy 0.46950 -15.52687 21.1 4.9 3.82 0.001
3 WLM Galaxy 0.47109 -15.50277 21.0 5.0 3.77 0.000
4 WLM Galaxy 0.47403 -15.52111 21.3 5.1 3.71 0.000
5 WLM Galaxy 0.47516 -15.53001 21.3 4.7 3.83 0.004
6 WLM Galaxy 0.47576 -15.50463 21.8 4.3 3.92 0.722
7 WLM Galaxy 0.47651 -15.50017 20.2 6.0 3.74 0.000
8 WLM Galaxy 0.47771 -15.51067 21.2 4.7 3.81 0.000
9 WLM Galaxy 0.47795 -15.49877 20.5 5.6 3.73 0.000
10 WLM Galaxy 0.47857 -15.52563 22.0 4.1 3.93 0.881
11 WLM Galaxy 0.48087 -15.49872 21.6 4.5 3.90 0.463
12 WLM Galaxy 0.48175 -15.49939 19.9 6.7 3.84 0.023
13 WLM Galaxy 0.48180 -15.50023 20.3 5.9 3.79 0.000
14 WLM Galaxy 0.48257 -15.50483 21.2 5.1 3.90 0.430
15 WLM Galaxy 0.48324 -15.51946 18.9 8.4 3.72 0.000
16 WLM Galaxy 0.48382 -15.50083 20.8 6.0 3.70 0.000
17 WLM Galaxy 0.48422 -15.52093 21.0 5.1 3.84 0.016
18 WLM Galaxy 0.48458 -15.50045 20.7 6.0 3.70 0.000
19 WLM Galaxy 0.48469 -15.50093 20.4 5.7 3.79 0.000
20 WLM Galaxy 0.48472 -15.53525 20.9 5.1 3.77 0.000
21 WLM Galaxy 0.48475 -15.49602 20.4 5.8 3.82 0.002
22 WLM Galaxy 0.48613 -15.51076 21.3 4.5 3.74 0.000
23 WLM Galaxy 0.48639 -15.50061 21.2 5.4 3.71 0.000
24 WLM Galaxy 0.48646 -15.52779 21.2 5.4 3.70 0.000
25 WLM Galaxy 0.48668 -15.49562 21.3 5.1 3.70 0.000
 [^13] Column 2: host galaxy. Columns 3–5: HST coordinates (deg) and magnitude (mag). Columns 6–7: MIST inferred stellar parameters. Column 8: Probability to be an extinguished MS star. The complete list will be made available in electronic form.
Table 4: Galaxies having more than 400 YSG candidates, sorted by distance.
Galaxy Distance HSTcov Database Camera / Filter Nsrc MIST M∗,ref $A_{V}$ [Fe/H] $n_{\mathrm{GM}}$ Ncand $f_{\mathrm{extin}}$ $f_{\mathrm{foreground}}$
@@
IC 10 0.7 49.1 HSCv3 ACS $F606W$ 208629 6 1.42 -1.30 10 485 0.214 0.237
M31 0.8 36.1 HSCv3+MAST ACS $F475W$ 31697166 4 0.49 -0.88 0 40566 0.260 0.128
M33 0.8 17.0 HSCv3 ACS $F475W$ 2834135 4 0.51 -1.13 10 11603 0.161 0.137
PegDIG 1.0 83.5 HSCv3 ACS $F475W$ 53724 3 0.28 -1.39 90 804 0.033 0.010
NGC 300 1.9 63.0 HSCv3 ACS $F606W$ 715947 4 0.05 -0.62 50 7525 0.316 0.013
NGC 1569 1.9 92.0 HSCv3 ACS $F606W$ 53637 6 0.81 -0.89 50 1019 0.437 0.005
NGC 55 2.2 15.0 HSCv3 ACS $F606W$ 146004 6 0.13 -1.45 10 1897 0.366 0.003
UGC 9240 2.8 100.0 HSCv3 ACS $F606W$ 25138 4 0.15 -1.13 10 832 0.239 0.002
NGC 2403 3.2 34.2 HSCv3 ACS $F606W$ 203205 8 0.27 -0.48 50 1494 0.391 0.016
NGC 2366 3.3 97.6 HSCv3 ACS $F555W$ 101143 6 0.20 -1.14 10 650 0.139 0.013
UGC 4305 3.4 58.7 HSCv3 ACS $F555W$ 97445 6 -0.04 -1.07 50 557 0.169 0.045
IC 342 3.4 6.9 HSCv3 ACS $F555W$ 12201 8 1.38 -0.87 0 449 0.152 0.202
NGC 7793 3.4 77.5 HSCv3 ACS $F555W$ 90987 8 0.06 -0.57 50 647 0.267 0.030
NGC 2976 3.8 67.0 HSCv3 ACS $F606W$ 100684 8 0.37 -0.71 50 571 0.460 0.003
M81 3.8 88.6 HSCv3 ACS $F606W$ 958597 8 0.20 -0.31 10 3357 0.344 0.043
NGC 253 3.9 38.8 HSCv3 ACS $F606W$ 634872 8 0.37 -0.93 10 4000 0.442 0.005
NGC 3077 3.9 55.4 HSCv3 ACS $F606W$ 121976 8 0.37 -0.67 50 495 0.297 0.006
NGC 4244 4.2 93.9 HSCv3 ACS $F606W$ 152981 8 0.12 -0.61 20 803 0.287 0.041
NGC 4449 4.3 84.4 HSCv3 ACS $F555W$ 142096 8 0.14 -0.79 20 2022 0.283 0.004
NGC 1313 4.3 69.7 HSCv3 ACS $F555W$ 168305 8 0.19 -0.80 0 1491 0.211 0.043
ESO 223-9 4.4 98.2 HSCv3 ACS $F606W$ 15109 8 0.60 -1.05 90 441 0.263 0.525
M94 4.6 34.3 HSCv3 ACS $F555W$ 17481 8 0.07 -0.35 50 496 0.300 0.035
ESO 137-18 4.7 91.9 HSCv3 ACS $F606W$ 19885 8 0.54 -0.83 30 441 0.176 0.452
M83 4.8 54.6 HSCv3 ACS $F555W$ 201354 8 0.16 -0.83 10 3671 0.349 0.031
NGC 4395 4.8 13.3 HSCv3 ACS $F606W$ 61908 6 0.10 -0.63 10 2604 0.146 0.009
NGC 4656 5.0 30.8 HSCv3 WFC3 $F555W$ 18113 8 -0.05 -0.48 50 554 0.316 0.009
NGC 6946 5.5 56.2 HSCv3 WFC3 $F606W$ 21289 8 0.97 -1.34 0 551 0.462 0.075
NGC 6503 6.2 61.5 HSCv3 ACS $F606W$ 39071 8 -0.02 -0.86 50 637 0.447 0.012
NGC 1566 6.5 41.6 HSCv3 ACS $F555W$ 21654 8 -0.01 -0.63 20 1004 0.231 0.012
Fourcade-Figueroa 6.7 43.3 HSCv3 ACS $F606W$ 38012 8 0.13 -1.26 0 446 0.308 0.137
M101 6.7 40.8 HSCv3 ACS $F555W$ 374087 8 0.12 -1.09 10 8885 0.278 0.018
IC 1727 6.9 59.4 HSCv3 ACS $F606W$ 21623 8 0.18 -0.97 40 608 0.220 0.022
NGC 3621 7.2 42.6 HSCv3 ACS $F555W$ 67015 8 0.17 -1.01 0 3089 0.398 0.011
M74 7.3 68.1 HSCv3 ACS $F555W$ 59490 8 0.38 -0.86 20 3265 0.267 0.007
IC 4710 7.4 97.8 HSCv3 ACS $F606W$ 26877 8 0.18 -0.73 20 569 0.280 0.127
M106 7.5 97.9 HSCv3 ACS $F555W$ 66349 10 0.32 -0.88 0 2173 0.244 0.029
NGC 4631 7.5 99.9 HSCv3 ACS $F606W$ 192224 10 0.17 -0.92 0 2496 0.269 0.003
NGC 2188 8.0 87.8 HSCv3 ACS $F606W$ 16356 8 0.20 -0.63 50 739 0.288 0.010
M51 8.6 50.9 HSCv3 ACS $F606W$ 58919 12 0.64 -0.46 0 576 0.295 0.016
NGC 4559 8.8 45.5 HSCv3 ACS $F606W$ 47602 10 0.14 -0.68 0 623 0.259 0.023
NGC 4485 8.9 99.5 HSCv3 ACS $F606W$ 10838 8 0.10 -0.65 10 456 0.340 0.004
NGC 2683 9.0 72.9 HSCv3 ACS $F606W$ 22402 8 -0.01 -0.63 20 472 0.361 0.011
NGC 24 9.1 86.9 HSCv3 ACS $F606W$ 18596 8 0.18 -1.26 0 415 0.402 0.010
NGC 4490 9.6 33.9 HSCv3 WFC3 $F555W$ 8312 8 0.15 -0.59 20 2847 0.146 0.000
NGC 3344 9.8 39.2 HSCv3 ACS $F606W$ 31474 10 0.37 -0.98 0 1378 0.247 0.009
NGC 672 10.0 57.5 HSCv3 ACS $F606W$ 32793 10 0.33 -1.18 0 984 0.322 0.012
IC 5201 10.3 62.0 HSCv3 ACS $F606W$ 22810 10 0.14 -0.48 50 463 0.235 0.045
M66 10.4 70.4 HSCv3 WFC3 $F555W$ 29283 12 0.15 -0.56 0 705 0.170 0.004
NGC 5398 11.4 97.8 HSCv3 ACS $F606W$ 14604 8 0.33 -1.11 0 552 0.388 0.028
NGC 4051 12.2 48.2 MAST WFC3 $F555W$ 5922 12 0.37 -0.68 50 1007 0.198 0.015
NGC 4565 12.8 73.5 HSCv3 ACS $F606W$ 42551 10 0.06 -0.24 80 926 0.313 0.028
@@
 [^14] Column 2: distance to the galaxy. Columns 3–6: information on their HST dataset. Columns 7–10: parameters used to fit their CMD. Column 11: number of resulting candidates. Columns 12–13: contamination estimates. The complete list of galaxies having YSG candidates will be made available in electronic form.